# Dynamic Tuning Towards Parameter and Inference Efficiency for ViT Adaptation

### v1

@@ 0.A.1 Details about Gumbel-Sigmoid
Gumbel-Softmax is proposed in [33] to conduct the differentiable sampling from a distribution. Given an unnormalized log probability  [^2] The original definition in [33] assumes the input $\{\mathbf{E}_{i}\}_{i=1}^{N}$ to be normalized log probability, while practical implementations [57] demonstrate its effectiveness even with unnormalized inputs. We also follow [57] to formulate and implement the Gumbel-Softmax. $\{\mathbf{E}_{i}\}_{i=1}^{N}$, the Gumbel-Softmax can be formulated as:
@@
obtaining the formulation of Gumbel-Sigmoid. Researchers in previous works, such as [22, 46], have also leveraged the Gumbel-Sigmoid formulation to facilitate end-to-end training of neural networks.
0.A.2 Implementation Details for each Task
@@
Following previous works [36, 37, 35], we fine-tune the model for 100 epochs on each dataset in VTAB-1K [74]. We do not use any data augmentation strategy in these experiments. We adopt the AdamW [50] optimizer. The learning rate is set to 1e-3 and gradually decays to 0 based on a cosine schedule [49].
@@
We adopt the settings in Table 7 to fine-tune the ViT with the proposed dynamic tuning. Experiments on other parameter-efficiency methods such as AdaptFormer [10], LoRA [32], and VPT [34] also follow the settings in Table 7. When we train a model with full tuning, we adopt a 1/10 base learning rate to make the training stable, otherwise, the model can not achieve reasonable results.
@@
  Configuration CIFAR-100 [39], SVHN [23], Food-101 [5]
Optimizer AdamW [50]
@@
Learning rate schedule Cosine decay [49]
@@
We adopt two video datasets, Kinetic-400 (K400) [9] and Something-Something V2 (SSv2) [24], to evaluate the performance as the token count scales up. The experimental settings are demonstrated in Table 8. Most of the settings are borrowed from [54]. The number of input frames is set to 8. We adopt multi-view during the test, which is a common practice in video action recognition. However, the original ViT lacks temporal modeling capabilities. To address this limitation, we draw inspiration from [71, 11]. By introducing a cross-attention layer after the ViT, along with a query token, we effectively aggregate temporal information across different frames. The final video action classification is performed based on this query token. Experiments on parameter-efficient fine-tuning methods also follow these settings. We adopt a 1/10 base learning rate for those experiments on full tuning.
Table 8: Experimental settings for video datasets. We follow most of settings in [54]. The number of input frames is set to 8 in all experiments. $lr=base\_lr\times batch\_size/256$
Configuration K400 [9] SSV2 [24]
Optimizer AdamW [50]
@@
Learning rate schedule Cosine decay [49]
@@
We conduct experiments on ADE20K [77] and COCO-stuff [6] to demonstrate the ability of DyT on dense prediction tasks. ADE20K contains 20,210 images from 150 fine-trained semantic concepts. COCO-Stuff consists of about 164k images with 172 semantic classes. The experimental settings are demonstrated in Table 9. All experiments for parameter-efficient fine-tuning methods also follow these settings. For the experiment on full tuning, we set adopt 1/10 learning rate for stable training and better performance.
@@
  Configuration ADE20K [77] COCO-stuff [6]
Optimizer AdamW [50]
@@
Learning rate schedule cosine decay [49]
@@
0.A.3 More Experiments on Image Datasets
To reflect the performance when the training data is sufficient, we further conduct experiments on 6 datasets including CIFAR-10 [39], DTD [13], Aircraft [51], Flowers [53], Pet [55], and Car [21]. We report the top1 accuracy for CIFAR-10, Car, and DTD, while providing the mean class accuracy for the remaining datasets. The experimental settings also follow the Table 7
We find that DyT achieves comparable performance to full tuning and AdaptFormer [10], while reducing the computational cost significantly. Notably, the performance of dynamic full tuning is unstable on some datasets, such as Aircraft and Car. This phenomenon may be attributed to the dynamic token mechanism, which can potentially impact pre-trained parameters when all parameters are trainable. This further verifies the importance of the proposed dynamic tuning.
Table 10: Results on image datasets. “Dynamic full tuning” denotes that we fine-tuning all parameters in the original ViT with the proposed token dispatcher. “Frozen” means that all parameters in the backbone keep frozen. “Param. (M)” denotes the number of trainable parameters in backbones. “FLOPSs (G)” is the average FLOPs on CIFAR-10. Bold font and underline denote the best and the second-best performance respectively.
CIFAR-10 DTD Aircraft Flowers Pet Car
Traditional methods
Full tuning 85.80 17.58 98.76 74.52 80.53 99.00 93.11 88.71
Linear 0 17.58 96.70 73.67 40.51 99.03 92.78 52.85
Dynamic full tuning 85.80 12.15 97.16 54.15 62.50 61.34 75.16 70.48
Parameter-efficient tuning methods
AdaptFormer [10] 1.19 17.81 99.13 77.61 78.23 99.49 94.46 87.33
LoRA [32] 1.19 17.58 98.88 75.53 76.32 99.50 93.62 87.00
VPT [34] 0.07 18.32 98.75 72.66 68.91 99.35 93.92 80.72
The proposed DyT
DyT 1.19 11.41 98.72 75.00 78.33 99.36 93.92 87.90
0.A.4 Generalization Capability of Transformer Architecture
To verify the generalization of the proposed dynamic tuning, we conduct experiments based on Swin-B [48]. The dynamic tuning can be directly applied to MLP blocks in the swin transformer without any modifications. The bottleneck dimension $d$ of the adapter in dynamic tuning also is set to 64. The results are demonstrated in Table 11. We can observe that the dynamic tuning can reduce both the tunable parameters and FLOPs while achieving comparable or even better performance across three datasets. This verifies the generalization capability of the dynamic tuning.
Table 11: Generalization Capability on Swin Transformer [48]. The experiments are conducted based on Swin-B. “Param. (M)” denotes the number of trainable parameters in backbones. “FLOPs (G)” denotes the average FLOPs on CIFAR-100. The bottleneck dimension $d$ is set to 64.
@@ 0.A.4 Generalization Capability of Transformer Architecture
0.A.5 Scaling-up Model Size
We verify the effectiveness of dynamic tuning when the model size is scaled up. We conduct experiments on ViT-L [19] and compare dynamic tuning with full tuning. The bottleneck dimension $d$ of the adapter in dynamic tuning is also set to 64. The results are demonstrated in Table 12. We can observe that with the activation rate set to 0.3, DyT has outperformed “full tuning” obviously on both CIFAR-100 and Food-101, while resulting in significantly lower computational cost.
We further compare the proposed dynamic tuning with full tuning on the VTAB-1K benchmark [74]. The results are demonstrated in Table 13. With only 0.44M tunable parameters and 43.52 GFLOPs, dynamic tuning surpasses full tuning across most datasets and achieves much better average performance.
Table 12: Scale up the model size to ViT-L [19]. “Param. (M)” denotes the number of trainable parameters in backbones. “FLOPs (G)” denotes the average FLOPs on CIFAR-100. The default setting is marked in color. The bottleneck dimension $d$ is set to 64.
@@ 0.A.5 Scaling-up Model Size
DyT $r=0.5$ 76.6 94.4 72.2 99.4 92.5 89.2 57.8 86.7 95.8 86.5 76.3 78.4 62.5 50.2 81.3 86.7 54.8 37.3 41.8 77.0 0.44 43.52
0.A.6 Investigation on the Temperature
We explore the temporal $\tau$ of the proposed dynamic tuning. The results are demonstrated in Table 14. When the temperature is smaller e.g. 0.1, the Gumbel Sigmoid tends to produce binary outputs close to 0 or 1. Conversely, larger temperatures lead to more uniform outputs, approaching 0.5. Results demonstrate that the performance is not too sensitive to the temperature and our model can achieve reasonable performance with all temperatures in the table. We also observe that the model with $\tau=1$ achieves the best performance on CIFAR-100, SVHN, and SSv2, while decaying the temperature with a schedule leads to the best result on Food-101 and K400, which shows that adjusting the temperature can help the model to achieve better performance. Since identifying the optimal temperature is not the primary focus of this paper, we directly set the temperature to 5 by default.
  Temperature Image Accuracy (%) $\uparrow$ Video Accuracy (%) $\uparrow$
CIFAR-100 SVHN Food-101 K400 SSv2
0.1 90.91 96.24 89.72 73.16 44.84
1 91.61 97.20 90.08 74.38 45.69
5 91.37 97.08 90.32 74.39 45.34
Schedule 91.58 97.13 90.39 74.57 45.51
Figure 8: Visualization of activated tokens. We present representative samples from the K400 [9] dataset. Blue patches represent the tokens activated in token dispatcher. Results verify that the token dispatcher has learned to identify informative tokens during fine-tuning. Zoom in for better view.
Figure 9: Visualization of activated tokens. We present representative samples from the SSv2 [24] dataset. Blue patches represent the tokens activated in token dispatcher. Results verify that the token dispatcher has learned to identify informative tokens during fine-tuning. Zoom in for better view.
0.A.7 Additional Visualizations of Activated Tokens
@@
We provide more visualizations of activated tokens from samples in K400 [9] and SSv2 [24] in Figure 8 and Figure 9, respectively. Results demonstrate that most activated tokens in higher layers e.g. Layer10 come from the primary objects. This proves that the proposed token dispatcher learns to activate informative tokens.

### v2

@@ A.7 Details about Gumbel-Sigmoid
Gumbel-Softmax is proposed in jang2016categorical to conduct the differentiable sampling from a distribution. Given an unnormalized log probability The original definition in jang2016categorical assumes the input $\{\mathbf{E}_{i}\}_{i=1}^{N}$ to be normalized log probability, while practical implementations pytroch_gumbel demonstrate its effectiveness even with unnormalized inputs. We also follow pytroch_gumbel to formulate and implement the Gumbel-Softmax. $\{\mathbf{E}_{i}\}_{i=1}^{N}$, the Gumbel-Softmax can be formulated as:
@@
obtaining the formulation of Gumbel-Sigmoid. Researchers in previous works, such as geng2020does ; liu2021gumbel , have also leveraged the Gumbel-Sigmoid formulation to facilitate end-to-end training of neural networks.
A.8 Implementation Details for each Task
@@
Following previous works jie2022convolutional ; jie2023fact ; jiang2023res , we fine-tune the model for 100 epochs on each dataset in VTAB-1K zhai2019large . We do not use any data augmentation strategy in these experiments. We adopt the AdamW loshchilov2017decoupled optimizer. The learning rate is set to 1e-3 and gradually decays to 0 based on a cosine schedule loshchilov2016sgdr .
@@
We adopt the settings in Table 14 to fine-tune the ViT with the proposed dynamic tuning. Experiments on other parameter-efficiency methods such as AdaptFormer chen2022adaptformer , LoRA hu2021lora , and VPT jia2022visual also follow the settings in Table 14. When we train a model with full tuning, we adopt a 1/10 base learning rate to make the training stable, otherwise, the model can not achieve reasonable results.
@@
  Optimizer AdamW loshchilov2017decoupled
@@
Learning rate schedule Cosine decay loshchilov2016sgdr
@@
We adopt two video datasets, Kinetic-400 (K400) carreira2017quo and Something-Something V2 (SSv2) goyal2017something , to evaluate the performance as the token count scales up. The experimental settings are demonstrated in Table 15. Most of the settings are borrowed from pan2022st . The number of input frames is set to 8. We adopt multi-view during the test, which is a common practice in video action recognition. However, the original ViT lacks temporal modeling capabilities. To address this limitation, we draw inspiration from yu2022coca ; chen2024context . By introducing a cross-attention layer after the ViT, along with a query token, we effectively aggregate temporal information across different frames. The final video action classification is performed based on this query token. Experiments on parameter-efficient fine-tuning methods also follow these settings. We adopt a 1/10 base learning rate for those experiments on full tuning.
Table 15: Experimental settings for video datasets. We follow most of settings in pan2022st . The number of input frames is set to 8 in all experiments. $lr=base\_lr\times batch\_size/256$
Configuration K400 carreira2017quo SSV2 goyal2017something
Optimizer AdamW loshchilov2017decoupled
@@
Learning rate schedule Cosine decay loshchilov2016sgdr
@@
We conduct experiments on ADE20K zhou2017scene and COCO-stuff caesar2018coco to demonstrate the ability of DyT on dense prediction tasks. ADE20K contains 20,210 images from 150 fine-trained semantic concepts. COCO-Stuff consists of about 164k images with 172 semantic classes. The experimental settings are demonstrated in Table 16. All experiments for parameter-efficient fine-tuning methods also follow these settings. For the experiment on full tuning, we set adopt 1/10 learning rate for stable training and better performance.
@@
  Configuration ADE20K zhou2017scene COCO-stuff caesar2018coco
Optimizer AdamW loshchilov2017decoupled
@@
Learning rate schedule cosine decay loshchilov2016sgdr
@@
A.9 Generalization Capability of Transformer Architecture
To verify the generalization of the proposed dynamic tuning, we conduct experiments based on Swin-B liu2021swin . The dynamic tuning can be directly applied to MLP blocks in the swin transformer without any modifications. The bottleneck dimension $d$ of the adapter in dynamic tuning also is set to 64. The results are demonstrated in Table 17. We can observe that the dynamic tuning can reduce both the tunable parameters and FLOPs while achieving comparable or even better performance across three datasets. This verifies the generalization capability of the dynamic tuning.
Table 17: Generalization Capability on Swin Transformer liu2021swin . The experiments are conducted based on Swin-B. “Param. (M)” denotes the number of trainable parameters in backbones. “FLOPs (G)” denotes the average FLOPs on CIFAR-100. The bottleneck dimension $d$ is set to 64.
@@ A.9 Generalization Capability of Transformer Architecture
A.10 Scaling-up Model Size
We verify the effectiveness of dynamic tuning when the model size is scaled up. We conduct experiments on ViT-L dosovitskiy2020image and compare dynamic tuning with full tuning. The bottleneck dimension $d$ of the adapter in dynamic tuning is also set to 64. The results are demonstrated in Table 18. We can observe that with the activation rate set to 0.3, DyT has outperformed “full tuning” obviously on both CIFAR-100 and Food-101, while resulting in significantly lower computational cost.
We further compare the proposed dynamic tuning with full tuning on the VTAB-1K benchmark zhai2019large . The results are demonstrated in Table 19. With only 0.44M tunable parameters and 43.52 GFLOPs, dynamic tuning surpasses full tuning across most datasets and achieves much better average performance.
Table 18: Scale up the model size to ViT-L dosovitskiy2020image . “Param. (M)” denotes the number of trainable parameters in backbones. “FLOPs (G)” denotes the average FLOPs on CIFAR-100. The default setting is marked in color. The bottleneck dimension $d$ is set to 64.
@@ A.10 Scaling-up Model Size
DyT $r=0.5$ 79.1 95.6 74.5 99.5 92.6 90.8 59.3 86.9 96.6 87.2 76.5 84.5 62.9 53.3 83.5 88.4 57.3 38.7 44.6 78.5 0.44 43.52
A.11 Additional Visualizations of Activated Tokens
@@
We provide more visualizations of activated tokens from samples in K400 carreira2017quo and SSv2 goyal2017something in Figure 8 and Figure 9, respectively. Results demonstrate that most activated tokens in higher layers e.g. Layer10 come from the primary objects. This proves that the proposed token dispatcher learns to activate informative tokens.
Figure 8: Visualization of activated tokens. We present representative samples from the K400 carreira2017quo dataset. Blue patches represent the tokens activated in token dispatcher. Results verify that the token dispatcher has learned to identify informative tokens during fine-tuning. Zoom in for better view.
Figure 9: Visualization of activated tokens. We present representative samples from the SSv2 goyal2017something dataset. Blue patches represent the tokens activated in token dispatcher. Results verify that the token dispatcher has learned to identify informative tokens during fine-tuning. Zoom in for better view.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Abstract
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Conclusion
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: See Method
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Experiments
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have released the code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Experiments
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: See Experiments
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Experiments
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: See the main paper
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: No negative societal impacts. Our model is only designed for fine-tuning.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Not release of data or models that have a high risk
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: we cite all related paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: See the main paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: See the main paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Not involve crowdsourcing nor research with human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.